OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (zarkon):

Use... \[\large\vec{a}\cdot\vec{b}=\|\vec{a}\|\|\vec{b}\|\cos(\theta)\]

OpenStudy (anonymous):

What's a and what's b?

OpenStudy (zarkon):

u and v if you want (any two vectors)

OpenStudy (zarkon):

theta is the angle between the vectors

OpenStudy (anonymous):

So it's -2*-5cos___

OpenStudy (zarkon):

?

OpenStudy (anonymous):

u*vcos(theta)

OpenStudy (zarkon):

\[\large\vec{u}\cdot\vec{v}=\|\vec{u}\|\|\vec{v}\|\cos(\theta)\]

OpenStudy (anonymous):

Ok, thank you!